# Association and Independence Test for Random Objects

### v2

@@ 1 Introduction
As data collected in scientific research have become increasingly complex, reaching beyond traditional scalar or vector formats, there is a need for statistical methodology and especially inference that can be applied for data located in more general spaces (Dryden and Kent, 2015; Dubey and Müller, 2020; Marron and Dryden, 2021; Dubey et al., 2024). Notable examples for such non-traditional and non-Euclidean data are distribution-valued data and Riemannian manifold-valued data, which have numerous applications in fields such as economics (Gunsilius, 2023), biology (Bolstad et al., 2003), brain imaging (Lila et al., 2016; Lila and Aston, 2020) and geography (Mitchell et al., 2025). For data residing in spaces that are endowed with a (pseudo-)Riemannian geometry, a common approach is to apply transformations, such as the Riemannian log map, to project the data into a linear tangent space (Petersen and Müller, 2016; Bigot et al., 2017; Lin and Yao, 2019; Chen et al., 2023). If such a linearizing map is available, one may adopt traditional methodologies for data in Euclidean spaces such as quantifications for association of multivariate or functional data with their usual Euclidean (or $\mathcal{L}^{2}$) metrics. However, the linearizing maps that are commonly employed are either metric-distorting or non-bijective, requiring ad hoc solutions to construct pseudo-inverses (Bigot et al., 2017; Chen et al., 2023). Also data may reside in metric spaces without an underlying Riemannian geometry, such as the space of phylogenetic trees (Billera et al., 2001).
@@
The study of data from general metric spaces has given rise to a prominent research area that has been referred to as metric statistics (Dubey et al., 2024), where random elements taking values in general metric spaces have been referred to as random objects (Müller, 2016; Bhattacharjee et al., 2025). The concept of a mean in general metric spaces was introduced by Fréchet (1948) and more recently extended to define conditional means in regression settings (Petersen and Müller, 2019) under the notion of Fréchet regression. These concepts are based on intrinsic metric functions and do not require any transformation to an extrinsic linear space. Such intrinsic approaches that directly confront the absence of linear operations have gained much interest in recent studies (Dubey and Müller, 2020; Zhu and Müller, 2023; Schötz, 2025; Jiang et al., 2024). Distance profiles (Dubey et al., 2024; Wang et al., 2024) corresponding to a class of one-dimensional distribution functions that characterize the distribution of random objects under mild conditions have proved instrumental for predictive conformal inference (Zhou and Müller, 2024a).
@@
In this paper, we consider the setting where $X$ and $Y$ are random objects that may be situated in different metric spaces and develop a unified framework to quantify the association between $X$ and $Y$. We extend the concept of distance profiles to random object pairs $(X,Y)$ and demonstrate that the joint distance profile function equals the product of the marginal distance profiles of $X$ and $Y$ if and only if $X$ and $Y$ are independent. Building on this basic result, we introduce the profile association, a dependence measure that is zero if and only if $X$ and $Y$ are independent and is one when $X$ and $Y$ have a specific functional relationship that we characterize below. The analysis of the empirical profile association relies on a connection between the Hoeffding D statistic and the proposed profile association, which makes it possible to devise a consistent permutation test for testing the independence between $X$ and $Y$ and to derive its power properties, the profile independence test. Popular alternatives to obtain inference with distance-based tests include the energy test and the d-variable Hilbert-Schmidt Independence Criterion (dHSIC), which have been extended to non-Euclidean data (Székely et al., 2007; Lyons, 2013; Sejdinovic et al., 2013) and we include these methods in our numerical comparisons.
For random elements taking values in Banach spaces, Pan et al. (2020) introduced an independence measure based on the concept of ball covariance, a concept closely related to distance profiles. However, the proof of the independence-zero equivalence property of ball covariance in Pan et al. (2020) depends on the linear structure of the Banach space. In practical applications, it is uncommon to encounter data that reside exclusively in a Banach space without also belonging to a Hilbert space equipped with an inner product, where one has available an array of methods that utilize the availability of linear operations (Lee et al., 2020; Lundborg et al., 2022). We emphasize that our proposed approach to quantify and infer dependence also applies to high-dimensional and functional data equipped with the Euclidean or $L^{2}$ metric, as it is universally applicable across all metric spaces.
@@
We note that recently Wang et al. (2024) extended the concept of ball covariance to metric spaces and this paper also features a subsection on its application for independence testing. However, since independence testing is only a secondary focus of this paper, the independence testing framework for random objects as presented remains incomplete, as the consistency properties along with the details and theoretical guarantees for the power of the resulting permutation implementation have not been established. Moreover, simulation studies demonstrate that our proposed profile independence test outperforms the ball covariance test across all settings. Very recently and independently of our work, another study on testing mutual independence between random objects was posted on arXiv (Chen and Dubey, 2025). While this work also utilizes distance profiles, it does not discuss the independence-zero equivalence property, and the proposed test statistic is based on estimated distance profiles. Our testing framework was derived simultaneously with this work, establishes the independence-zero equivalence, provides details for the dependence-1 equivalence, is based on a different Hoeffding D-statistic-based test statistic, employs different proof techniques and utilizes a half-permutation test for the study of consistency and power. Furthermore, we extend our approach to the important case of testing and quantifying conditional associations, when conditioning on a Euclidean random vector.
@@
The quantification of conditional associations and inference is motivated by the fact that many real-world datasets involve complex random objects along with Euclidean predictors, such as age, time or experimental conditions. Understanding conditional relationships then enables researchers to model, predict, and interpret such data more effectively, especially in fields such as medical imaging, neuroscience, genetics, and environmental science. Specifically, we introduce the notion of conditional profile association to quantify the dependence between random objects $X$ and $Y$ given a Euclidean predictor $Z$ We showcase the application of conditional profile association for human longevity, where we study the relationship between the age-at-death distributions of males and females across different calendar years for various countries in Section 6.2. Using estimated joint and marginal distance profile functions, we also propose a statistic to carry out a conditional profile independence test to infer whether $X$ and $Y$ are conditionally independent and establish the uniform convergence of the estimated joint and marginal conditional distance profiles, which then implies the uniform convergence of conditional profile association. We also study the asymptotic distribution of the test statistic for the test for conditional independence under the null hypothesis and its implementation with a permutation test.
@@ 2 Profile association for random objects
We state here some basic assumptions and notations. In the following, $X$ and $Y$ are random elements that take values in totally bounded, separable metric spaces $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ with associated metrics $d_{1}(\cdot,\cdot)$ and $d_{2}(\cdot,\cdot)$, respectively. We refer to such random elements as random objects. We denote by $(S,\mathcal{S},\mathbb{P})$ an underlying probability space, where $\mathcal{S}$ is the Borel $\sigma$-algebra on $S$ and $\mathbb{P}$ is a probability measure, and by $\mu_{XY}$ the joint measure of $X$ and $Y$, where $\mu_{XY}(A\times B)=\mathbb{P}(s\in S:X(s)\in A\text{ and }Y(s)\in B)$ for any Borel measurable sets $A\subset\mathcal{M}_{1}$ and $B\subset\mathcal{M}_{2}$. We further assume that the marginal distributions of $X$ and $Y$, denoted by $\mu_{X}$ and $\mu_{Y}$, have finite first-order moments, i.e., $\int d_{1}(\omega_{1},x)\,\mathrm{d}\mu_{X}(x)<\infty$ and $\int d_{2}(\omega_{2},x)\,\mathrm{d}\mu_{Y}(x)<\infty$ for some (and thus any) $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$.
We are interested in quantifying the association between $X$ and $Y$, and especially in determining whether $X$ and $Y$ are independent, which is equivalent to $\mu_{XY}=\mu_{X}\times\mu_{Y}$. The following proposition states a necessary and sufficient condition for $X$ and $Y$ to be independent. In general metric spaces, linear operations like addition, subtraction and inner product that are fundamental to conventional statistical inference for Euclidean data are not available. Furthermore, in some metric spaces like $\mathcal{L}^{2}{[0,1]}$, the absence of a well-defined Lebesgue measure renders density-based inference methods infeasible. Since the only structure available in general metric spaces is the distance between two elements, developing intrinsic statistical methodologies that align with the underlying geometry requires leveraging these distances. The following proposition provides a sufficient condition under which the independence of probability measures is equivalent to the independence of distances between random elements and every point in the metric space.
@@
If the probability measures $\mu_{X}$, $\mu_{Y}$, and $\mu_{XY}$ are fully determined by their values on closed balls, then $\mu_{XY}=\mu_{X}\times\mu_{Y}$ if and only if $d_{1}(X,\omega_{1})$ is independent of $d_{2}(Y,\omega_{2})$ for all $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$.
A probability measure $\mu$ is said to be “fully determined by its values on closed balls” if knowing the values $\mu(\overline{\mathcal{B}}(\omega,r))$ for all points $\omega\in\mathcal{M}$ and all radii $r>0$ is sufficient to uniquely specify $\mu$ on all Borel sets. This problem was first proposed and studied by Christensen (1970), Davies (1971), and Hoffmann-Jørgensen (1975), and has been recently investigated by Blanche et al. (2016), along with applications in statistics and probability discussed in Lyons (2013), Dubey et al. (2024), and Chen and Dubey (2025). The following proposition provides two sufficient conditions under which the condition in Proposition 1 holds. A probability measure $\mu$ on a metric space $(\mathcal{M},d)$ is said to be doubling if there exists a constant $L_{\mu}\geq 0$ such that
where ¯$\overline{\mathcal{B}}(\omega,r):=\left\{\omega^{\prime}\in\mathcal{M}:d(\omega,\omega^{\prime})\leq r\right\}$ is the closed ball centered at $\omega$ with radius $r$, and $\mathrm{supp}(\mu):=\left\{\omega^{\prime}\in\mathcal{M}:\mu(\{\omega^{\prime}\})>0\right\}$ is the support of $\mu$. The concept of a doubling measure in metric spaces was first introduced by Beurling and Ahlfors (1956) and has since become a fundamental notion in analysis on metric spaces, including various areas such as harmonic analysis and geometric measure theory (Heinonen, 2001).
Proposition 2.
Proposition 1 holds provided that either of the following assumptions is satisfied:
@@
    The measures $\mu_{X}$, $\mu_{Y}$ and $\mu_{XY}$ satisfy the doubling condition (1).
Assumption M0 pertains to the structural properties of the underlying metric spaces. The proof is based on Proposition 3.5, Proposition 3.7, and Lemma 3.8 in Lyons (2013). Preliminary background on metric spaces of strong negative type is provided in Appendix A, with additional details and illustrative examples available in Klebanov et al. (2005) and Lyons (2013). Assumption P0 concerns the properties of the associated probability measures and was employed in recent work Chen and Dubey (2025) to study mutual independence in metric spaces. Proposition 1 and Proposition 2 establish a connection between the independence of probability measures and the independence of distances of random elements, providing the basis for statistical inference for association and independence in metric spaces.
Our approach utilizes distance profiles (Dubey et al., 2024), where for each $\omega_{1}\in\mathcal{M}_{1}$, the marginal distance profile with respect to $\mu_{X}$ is defined as
$F^{X}_{\omega_{1}}(u)=\mathbb{P}\left(d_{1}(X,\omega_{1})\leq u\right),\,\,\,u\geq 0,\,\,\omega_{1}\in\mathcal{M}_{1},$ (2)
and marginal distance profiles $F^{Y}_{\omega_{2}}(u)$ for each $\omega_{2}\in\mathcal{M}_{2}$ with respect to $\mu_{Y}$ are defined analogously. Distance profiles establish an intrinsic link between random objects and the one-dimensional distribution functions $F_{\omega}$ and provide a feasible framework for developing inference tools for random objects. This framework has been previously applied for the construction of two-sample tests (Dubey et al., 2024; Wang et al., 2024) and conformal prediction (Zhou and Müller, 2024a).
To measure the independence between $X$ and $Y$, we adopt the idea to extend (2) to the two-dimensional case by considering joint distance profiles for pairs of elements $(\omega_{1},\omega_{2})$,   $\omega_{1}\in\mathcal{M}_{1},\omega_{2}\in\mathcal{M}_{2}$, with respect to $\mu_{XY}$,
$F_{\omega_{1},\omega_{2}}^{XY}(u,v)=\mathbb{P}\left(d_{1}(X,\omega_{1})\leq u,\,d_{1}(Y,\omega_{2})\leq v\right),\quad\,u,v\geq 0,\,\,\,\omega_{1}\in\mathcal{M}_{1},\,\,\omega_{2}\in\mathcal{M}_{2}.$ (3)
Proposition 2 then implies the following result, which relates the independence between random objects to their corresponding distance profiles.
Proposition 3.
Under either (M0) or (P0), the random objects $X$ and $Y$ are independent if and only if
$F_{\omega_{1},\omega_{2}}^{XY}(t,s)=F_{\omega_{1}}^{X}(t)F_{\omega_{2}}^{Y}(s),$
for all $t,s\geq 0$, $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$.
Proposition 3 enables us to transform the independence testing problem for random objects into testing a class of two-dimensional vectors $\{(d_{1}(X,\omega_{1}),d_{2}(Y,\omega_{2}))^{\top}\}$ indexed by $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$. Extensive research has been conducted on testing the independence of vector-valued random variables $X$ and $Y$ by quantifying the discrepancy between their joint measure and the product of their marginal measures. For scalar random variables $X$ and $Y$ with joint distribution function $F(x,y)$, in a seminal paper Hoeffding (1948) introduced the functional
@@
which corresponds to a type of profile association between $X$ and $Y$. When $F(x,y)$ has a continuous density function, $X$ and $Y$ are independent if and only if $\Delta(F)=0$.
@@
where $(X^{\prime},Y^{\prime})$ is an independent copy of $(X,Y)$, and the random distance profiles $F_{X^{\prime},Y^{\prime}}^{XY}(u,v)$, $F_{X^{\prime}}^{X}(u)$, and $F_{Y^{\prime}}^{Y}(v)$ are defined as:
$F_{X^{\prime},Y^{\prime}}^{XY}(u,v):=\mathbb{P}_{X,Y}\left(d_{1}(X,X^{\prime})\leq u,\,d_{2}(Y,Y^{\prime})\leq v\right),$
$F_{X^{\prime}}^{X}(u):=\mathbb{P}_{X}\left(d_{1}(X,X^{\prime})\leq u\right),\quad F_{Y^{\prime}}^{Y}(v):=\mathbb{P}_{Y}\left(d_{2}(Y,Y^{\prime})\leq v\right),$
for all $u,v\geq 0$. Throughout this paper, the subscript in $\mathbb{E}_{X}$ (or $\mathbb{P}_{X}$) indicates that the expectation (or probability) is taken over the random variable(s) specified in the subscript only. For a random object $X$, i.e., a metric space-valued random variable, we denote by $X(s)\in\mathcal{M}$ the value it assumes in the metric space $\mathcal{M}$ for a specific argument $s$ that is an element of the underlying probability space $S$.
@@
The assumption in Theorem 1 is relatively mild. Specifically, when $\mathcal{M}_{1}=\mathcal{M}_{2}=\mathbb{R}$, this assumption can be derived from the condition in Hoeffding (1948), which requires that the pair $(X,Y)$ has a continuous joint density function. The first statement of Theorem 1 shows that the proposed profile association measure satisfies both I-consistency and D-consistency; see Weihs et al. (2018):
  • I-consistency. If $X$ and $Y$ are independent, the profile association measure is zero.
  • D-consistency. If $X$ and $Y$ are dependent, the profile association measure is nonzero.
A similar independence-zero equivalence property has been established for ball covariance (Pan et al., 2020), which was proposed to test independence between Banach space-valued random elements (while we consider here random objects which take values in more general metric spaces). The proof of Theorem 1 relies on a mapping between negative-type metric spaces and Hilbert spaces (Schoenberg, 1937, 1938; Lyons, 2013).
@@
The proposed profile association measure (4) specifically targets the quantification of association for random objects in general metric spaces. We note that monotone functions are not well-defined in metric spaces that lack a natural order and furthermore that alternative measures of independence often assign the same maximal association level to any monotone (or measurable) function of $X$, offering no additional insight into the degree of dependence between $Y$ and $X$. This limitation prevents these measures from distinguishing varying levels of dependence. In contrast, the proposed profile association uniquely identifies $Y=X$ almost surely as the maximal form of dependence, providing a more precise interpretation of maximal dependence. This property also indicates that test statistics based on (4) may have greater power in detecting dependence, and we find that this is indeed the case in the simulations reported in Section 5.

### v3

@@ 1 Introduction
As data collected in scientific research have become increasingly complex, reaching beyond traditional scalar or vector formats, there is a need for statistical methodology and especially inference that can be applied for data located in more general spaces (Dryden and Kent, 2015; Dubey and Müller, 2020; Marron and Dryden, 2021; Dubey et al., 2024). Notable examples for such non-traditional and non-Euclidean data are distribution-valued data and Riemannian manifold-valued data, which have numerous applications in fields such as economics (Gunsilius, 2023), biology (Bolstad et al., 2003), brain imaging (Lila et al., 2016; Lila and Aston, 2020) and geography (Mitchell et al., 2025). For data residing in spaces that are endowed with a (pseudo-)Riemannian geometry (O’Neill, 1983), a common approach is to apply transformations, such as the Riemannian log map, to project the data into a linear tangent space (Petersen and Müller, 2016; Bigot et al., 2017; Lin and Yao, 2019; Chen et al., 2023). If such a linearizing map is available, one may adopt traditional methodologies for data in Euclidean spaces such as quantifications for association of multivariate or functional data with their usual Euclidean (or $\mathcal{L}^{2}$) metrics. However, the linearizing maps that are commonly employed are either metric-distorting or non-bijective, requiring ad hoc solutions to construct pseudo-inverses (Bigot et al., 2017; Chen et al., 2023). Also data may reside in metric spaces without an underlying Riemannian geometry, such as the space of phylogenetic trees (Billera et al., 2001).
@@
The study of data from general metric spaces has given rise to a prominent research area that has been referred to as metric statistics (Dubey et al., 2024), where random elements taking values in general metric spaces have been referred to as random objects (Müller, 2016; Bhattacharjee et al., 2025). The concept of a mean in general metric spaces was introduced by Fréchet (1948) and more recently extended to define conditional means in regression settings (Petersen and Müller, 2019) under the notion of Fréchet regression. These concepts are based on intrinsic metric functions and do not require any transformation to an extrinsic linear space. Such intrinsic approaches that directly confront the absence of linear operations have gained much interest in recent studies (Dubey and Müller, 2020; Zhu and Müller, 2023; Schötz, 2025; Jiang et al., 2024). Distance profiles (Dubey et al., 2024; Wang et al., 2024) corresponding to a class of one-dimensional distribution functions that characterize the distribution of random objects under mild conditions have proved instrumental for predictive conformal inference (Zhou and Müller, 2025).
@@
In this paper, we consider the setting where $X$ and $Y$ are random objects that may be situated in different metric spaces and develop a unified framework to quantify the association between $X$ and $Y$. We extend the concept of distance profiles to random object pairs $(X,Y)$ and demonstrate that the joint distance profile function equals the product of the marginal distance profiles of $X$ and $Y$ if and only if $X$ and $Y$ are independent. Building on this basic result, we introduce the profile association, a dependence measure that is zero if and only if $X$ and $Y$ are independent and is one when $X$ and $Y$ have a specific functional relationship that we characterize below. The analysis of the empirical profile association relies on a connection between the Hoeffding D statistic and the proposed profile association, which makes it possible to devise a consistent permutation test for testing the independence between $X$ and $Y$ and to derive its power properties, the profile independence test. Popular alternatives to obtain inference with distance-based tests include the energy test and ball divergence test, which have been extended to non-Euclidean data (Székely et al., 2007; Lyons, 2013; Sejdinovic et al., 2013; Wang et al., 2024; Gao et al., 2025) and we include these methods in our numerical comparisons.
Testing independence between non-Euclidean data has become an active area of research with important applications across many interdisciplinary fields. For random elements taking values in Banach spaces, Pan et al. (2020) introduced an independence measure based on the concept of ball covariance, which is closely related to distance profiles. However, their proof of the independence-zero equivalence relies on the linear structure of Banach spaces. In practice, it is uncommon to encounter data that reside exclusively in Banach spaces without also belonging to a Hilbert space, where one can leverage inner product structures and apply a wide range of linear methods (Lee et al., 2020; Lundborg et al., 2022). Recently, Wang et al. (2024) extended the notion of ball covariance to general metric spaces and included a subsection on its application to independence testing, along with an analysis of the asymptotic distribution of the test statistic under the null hypothesis. However, the overall framework for independence testing with random objects remains incomplete: key properties such as the consistency of the test and the power analysis of the permutation-based implementation have not yet been fully established.
We systematically study the independence testing problem for random objects based on distance profiles. Intrinsically, both the metric distribution functions (MDF) Wang et al. (2024) and the distance profiles Dubey et al. (2024) rely on distribution functions that characterize probabilistic information about distances in a metric space. In the context of independence testing, there are at least three advantages of using distance profiles over MDF. First, the framework of distance profiles naturally connects the independence of random objects with the independence of a class of one-dimensional random variables and satisfies a factorization property (see Proposition 1 in Section 2). This connection bridges the gap between dependence structures of complex objects and those of univariate random variables, allowing one to apply a wide range of existing tests for univariate independence. Second, because distance profiles are univariate distribution functions, it is straightforward to relate them to Hoeffding’s $D$ statistic (Hoeffding, 1948), enabling the construction of interpretable and computationally efficient test statistics. Third, the assumptions required for MDF and distance profiles to fully characterize probability measures on metric spaces are different. In Wang et al. (2024), it relies on the directionally $(\epsilon,\eta,L)$-limited condition (Federer, 2014), which is somewhat abstract and often difficult to verify, particularly in infinite-dimensional spaces such as Wasserstein spaces. In contrast, under the distance profile framework, this one-to-one correspondence holds under mild assumptions on either the metric space or the probability measure, making the approach more flexible and broadly applicable.
@@
Very recently and independently of our work, another study on testing mutual independence between random objects was posted on arXiv (Chen and Dubey, 2025). While this work also utilizes distance profiles, its proposed test statistic is based on estimated distance profiles. Our testing framework, developed simultaneously but independently, differs in several key aspects. It is based on a Hoeffding’s $D$-statistic-type test, establishes the dependence-1 equivalence, employs different proof techniques and employs a half permutation test to study consistency and power. Furthermore, we extend our approach to the important problem of testing and quantifying conditional associations, conditioning on a Euclidean random vector.
@@
The quantification of conditional associations and inference is motivated by the fact that many real-world datasets involve complex random objects along with Euclidean predictors, such as age, time or experimental conditions. Understanding conditional relationships enables researchers to model, predict, and interpret such data effectively, especially in fields such as medical imaging, neuroscience, genetics or environmental science. Specifically, we introduce the notion of conditional profile association to quantify the dependence between random objects $X$ and $Y$ given a Euclidean predictor $Z$. Using estimated joint and marginal distance profile functions, we also propose a statistic to carry out a conditional profile independence test to infer whether $X$ and $Y$ are conditionally independent and establish the uniform convergence of the estimated joint and marginal conditional distance profiles, which then implies the uniform convergence of conditional profile association. We also study the asymptotic distribution of the test statistic for the test for conditional independence under the null hypothesis and its implementation with a permutation test. We evaluate the finite-sample performance of the proposed conditional test through simulation studies and an application to human mortality data, where we examine conditional profile associations between male and female age-at-death distributions across different calendar years (see Section 6.2).
@@ 2 Profile association for random objects
We state here basic assumptions and notations. In the following, $X$ and $Y$ are random elements that take values in totally bounded separable metric spaces $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ with metrics $d_{1}(\cdot,\cdot)$ and $d_{2}(\cdot,\cdot)$, respectively. We refer to such random elements as random objects. We denote by $(S,\mathcal{S},\mathbb{P})$ an underlying probability space, where $\mathcal{S}$ is the Borel $\sigma$-algebra on $S$ and $\mathbb{P}$ is a probability measure, and by $\mu_{XY}$ the joint measure of $X$ and $Y$, where $\mu_{XY}(A\times B)=\mathbb{P}(s\in S:X(s)\in A\text{ and }Y(s)\in B)$ for any Borel measurable sets $A\subset\mathcal{M}_{1}$ and $B\subset\mathcal{M}_{2}$. The marginal distributions of $X$ and $Y$, denoted by $\mu_{X}$ and $\mu_{Y}$, are assumed to have finite first-order moments, i.e., $\int d_{1}(\omega_{1},x)\,\mathrm{d}\mu_{X}(x)<\infty$ and $\int d_{2}(\omega_{2},x)\,\mathrm{d}\mu_{Y}(x)<\infty$ for some (and thus any) $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$. We aim to quantify the association between $X$ and $Y$ and to determine whether $X$ and $Y$ are independent, which is equivalent to $\mu_{XY}=\mu_{X}\times\mu_{Y}$. In general metric spaces, linear operations like addition, subtraction and inner product that are fundamental to conventional statistical inference for Euclidean data are not available. Furthermore, in some metric spaces like $\mathcal{L}^{2}{[0,1]}$, the absence of a well-defined Lebesgue measure renders density-based inference methods infeasible. Since the only structure available in general metric spaces is the distance between two elements, developing intrinsic statistical methodologies that align with the underlying geometry requires working directly with distances.
We further note that to develop statistical methods for general metric spaces, one needs to restrict attention to a subclass of probability measures that are fully determined by their values on closed balls. Specifically, this means that the values $\mu(\overline{\mathcal{B}}(\omega,r))$ for all points $\omega\in\mathcal{M}$ and all radii $r>0$ uniquely determines the measure $\mu$ on all Borel sets, where $\overline{\mathcal{B}}(\omega,r):=\left\{\omega^{\prime}\in\mathcal{M}:d(\omega,\omega^{\prime})\leq r\right\}$ denotes the closed ball centered at $\omega$ with radius $r$. This property is commonly referred to as the one-to-one correspondence and was first studied by Christensen (1970), Davies (1971), and Hoffmann-Jørgensen (1975). More recently, it has been investigated by Blanche et al. (2016), with related applications in statistics and probability discussed in Lyons (2013), Dubey et al. (2024), and Chen and Dubey (2025).
Distance profiles (Dubey et al., 2024) provide a natural framework for studying the one-to-one correspondence and for developing inference tools for random objects by establishing an intrinsic connection between such objects and a class of one-dimensional distribution functions. This general framework has recently been applied for the construction of two-sample tests (Dubey et al., 2024; Wang et al., 2024) and for conformal prediction (Zhou and Müller, 2025). Specifically, for each $\omega_{1}\in\mathcal{M}_{1}$, the marginal distance profile with respect to $\mu_{X}$ is defined as
$F^{X}_{\omega_{1}}(u)=\mathbb{P}\left(d_{1}(X,\omega_{1})\leq u\right),\,\,\,u\geq 0,$ (1)
and the marginal distance profiles $F^{Y}_{\omega_{2}}(u)$ for each $\omega_{2}\in\mathcal{M}_{2}$ with respect to $\mu_{Y}$ are defined analogously. To quantify the dependence and characterize independence between $X$ and $Y$, we extend the marginal distance profiles in (1) to a two-dimensional version by introducing joint distance profiles for each pair $(\omega_{1},\omega_{2})\in\mathcal{M}_{1}\times\mathcal{M}_{2}$ with respect to $\mu_{XY}$:
$F_{\omega_{1},\omega_{2}}^{XY}(u,v)=\mathbb{P}\left(d_{1}(X,\omega_{1})\leq u,\,d_{2}(Y,\omega_{2})\leq v\right),\quad\,u,v\geq 0.$ (2)
The concept of joint distance profiles is a natural extension of distance profiles and has also been proposed independently in Chen and Dubey (2025). The following factorization proposition is a fundamental result that enables us to transform the independence testing problem for random objects into testing a class of two-dimensional vectors.
@@
Suppose that either of the following assumptions holds:
@@
    The measures $\mu_{X}$, $\mu_{Y}$, and $\mu_{XY}$ are assumed to satisfy the doubling condition:
    for $\mu=\mu_{X},\mu_{Y},$ and $\mu_{XY}$, where $\mathrm{supp}(\mu):=\left\{\omega^{\prime}\in\mathcal{M}:\mu(\{\omega^{\prime}\})>0\right\}$ denotes the support of $\mu$ and $L_{\mu}$ is a positive constant.
Then the random objects $X$ and $Y$ are independent if and only if
$F_{\omega_{1},\omega_{2}}^{XY}(t,s)=F_{\omega_{1}}^{X}(t)\cdot F_{\omega_{2}}^{Y}(s),\quad\text{for all }t,s\geq 0,\ \omega_{1}\in\mathcal{M}_{1},\ \omega_{2}\in\mathcal{M}_{2}.$
Assumption M0 relates to the structural properties of the underlying metric spaces. Under this assumption, Corollary 3.20 of Lyons (2013) shows that the product space $\mathcal{M}_{1}\times\mathcal{M}_{2}$ is also of strong negative type. A brief review of the strong negative type assumption is provided in Supplement, with additional details and examples available in Klebanov et al. (2005) and Lyons (2013). Assumption P0 imposes conditions on the probability measures and was adopted in recent work by Chen and Dubey (2025) to study mutual independence in metric spaces. The concept of a doubling measure in metric spaces was first introduced by Beurling and Ahlfors (1956) and has since become a fundamental notion for metric measure spaces, harmonic analysis and geometric measure theory (Heinonen, 2001). The significance of Proposition 1 is that it reduces the complex independence testing problem for random objects to the simpler problem of studying the dependence of two-dimensional vectors $\{(d_{1}(X,\omega_{1}),d_{2}(Y,\omega_{2}))^{\top}\}$, indexed by $\omega_{1}\in\mathcal{M}_{1}$ and $\omega_{2}\in\mathcal{M}_{2}$, where one can draw on extensive existing research.
A standard approach for testing the independence of vector-valued random variables $X$ and $Y$ is to quantify the discrepancy between their joint measure and the product of their marginal measures. For scalar random variables $X$ and $Y$ with joint distribution function $F(x,y)$, in a seminal paper Hoeffding (1948) introduced the functional
@@
which quantifies the association between $X$ and $Y$. When $F(x,y)$ has a continuous density function, $X$ and $Y$ are independent if and only if $\Delta(F)=0$.
@@
where $(X^{\prime},Y^{\prime})$ is an independent copy of $(X,Y)$, and the random distance profiles $F_{X^{\prime},Y^{\prime}}^{XY}(u,v)$, $F_{X^{\prime}}^{X}(u)$, and $F_{Y^{\prime}}^{Y}(v)$ are defined as
$F_{X^{\prime},Y^{\prime}}^{XY}(u,v)=\mathbb{P}_{X,Y}\left(d_{1}(X,X^{\prime})\leq u,\,d_{2}(Y,Y^{\prime})\leq v\right),$
$F_{X^{\prime}}^{X}(u)=\mathbb{P}_{X}\left(d_{1}(X,X^{\prime})\leq u\right),\quad F_{Y^{\prime}}^{Y}(v)=\mathbb{P}_{Y}\left(d_{2}(Y,Y^{\prime})\leq v\right),$
for all $u,v\geq 0$. Throughout this paper, the subscript in $\mathbb{E}_{X}$ (or $\mathbb{P}_{X}$) indicates that the expectation (or probability) is taken over the random variable(s) specified in the subscript only. For a random object $X$, i.e., a metric space-valued random variable, we denote by $X(s)\in\mathcal{M}_{1}$ the value it assumes in the metric space $\mathcal{M}_{1}$ for a specific argument $s$, which is an element of the underlying probability space $S$.
@@
The assumption in Theorem 1 is relatively mild. Specifically, when $\mathcal{M}_{1}=\mathcal{M}_{2}=\mathbb{R}$, this assumption can be derived from the condition in Hoeffding (1948), which requires that the pair $(X,Y)$ has a continuous joint density function. The first statement of Theorem 1 shows that the proposed profile association measure satisfies both I-consistency and D-consistency (Weihs et al., 2018):
  • I-consistency.
    If $X$ and $Y$ are independent, the profile association measure is zero.
  • D-consistency.
    If $X$ and $Y$ are dependent, the profile association measure is nonzero.
A similar independence-zero equivalence property has been established for ball covariance (Pan et al., 2020), which was proposed to test independence for the special case of Banach (and thus vector-)space-valued random elements. The proof of Theorem 1 utlizes a well-known mapping between negative-type metric spaces and Hilbert spaces (Schoenberg, 1937, 1938; Lyons, 2013).
@@
The proposed profile association measure (4) specifically targets the quantification of association for random objects in general metric spaces. We note that monotone functions are not well-defined in metric spaces that lack a natural order and furthermore that alternative measures of independence often assign the same maximal association level to any monotone (or measurable) function of $X$, providing no additional insight into the degree of dependence between $Y$ and $X$. This limitation prevents these measures from distinguishing varying levels of dependence. In contrast, the proposed profile association uniquely identifies $Y=X$ almost surely as the maximal form of dependence, providing a more precise interpretation of maximal dependence. This property also indicates that test statistics based on (4) may have greater power in detecting dependence, and we find that this is indeed the case in the simulations reported in Section 5.